Grade 12th passAIPMT

A capacitor of 1microF can withstand a maximum voltage of 6KV while another of 2 micro F can withstand max voltage of 4KV.If two are connected in series they withstand maximum voltage of what.?

Parul Rajput
9 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
Dear student,

Maximum charge that the first capacitor can store=CV=1\mu F\times 6kV=6mC
Maximum charge that the second capcitor can store=2\mu F\times 4kV=8mC

Hence charge more than6mCcannot be flown through the system.
Equivalent capapcitance of the system=\dfrac{2}{3}\mu F

Hence the breakdown voltage of the system=\dfrac{6mC}{\dfrac{2}{3}\mu F}=9kV